AAVSOnet status, 2016-05-12

Thu, 05/12/2016 - 10:44

Here is the current operating status of the network telescopes.

BSM_South  working nominally.

BSM_Berry  working nominally.  Greg Bolt replaced the control board in the CGEM mount, solving its problems.

BSM_HQ  working nominally, though we need an alternative to astrometry.net for initial pointing.

BSM_Hamren  down until the local computer is replaced.  The telescope is also getting moved to a new location soon and there will be a few nights of setup required.  Estimate is two weeks before data collection resumes.

BSM_NM  still trying to understand its new mount/software.  We're hoping to have this solved in a couple of days.

coker30 working nominally.

SRO working nominally.

OC61 working nominally.

TMO61  equipment working, but need to get back into the routine of observing.  We will be talking with the site manager in the next few days.